The sum of three integers is 92. The second number is three times the first number. The third number is ten less than twice the first number. Find the integers.


Answer by ikleyn(52788)

.

x + 3x + (2x-10) = 92


6x = 92 + 10


6x = 102


x = 102/6 = 17


etc.